# KaPO: Knowledge-aware Preference Optimization for Controllable Knowledge Selection in Retrieval-Augmented Language Models

### v1

@@ Abstract
By integrating external knowledge, Retrieval-Augmented Generation (RAG) has become an effective strategy for mitigating the hallucination problems that large language models (LLMs) encounter when dealing with knowledge-intensive tasks. However, in the process of integrating external non-parametric supporting evidence with internal parametric knowledge, inevitable knowledge conflicts may arise, leading to confusion in the model’s responses. To enhance the knowledge selection of LLMs in various contexts, some research has focused on refining their behavior patterns through instruction-tuning. Nonetheless, due to the absence of explicit negative signals and comparative objectives, models fine-tuned in this manner may still exhibit undesirable behaviors in the intricate and realistic retrieval scenarios. To this end, we propose a Knowledge-aware Preference Optimization, dubbed KaPO, aimed at achieving controllable knowledge selection in real retrieval scenarios. Concretely, we explore and simulate error types across diverse context combinations and learn how to avoid these negative signals through preference optimization methods. Simultaneously, by adjusting the balance between response length and the proportion of preference data representing different behavior patterns, we enhance the adherence capabilities and noise robustness of LLMs in a balanced manner. Experimental results show that KaPO outperforms previous methods for handling knowledge conflicts by over 37%, while also exhibiting robust generalization across various out-of-distribution datasets.
@@ I. Introduction
Large Language Models (LLMs) (Taylor et al. 2022; Zhao et al. 2023b) have been widely applied in various fields, such as natural language processing, question-answering systems, and text generation, giving rise to numerous AI applications (Kaplan et al. 2020; Vu et al. 2024). These models exhibit outstanding performance in many tasks, primarily due to their large-scale parameters and extensive pre-training data (Ziegler et al. 2020; Wang et al. 2023d). However, because of the static nature of the training data, LLMs may generate seemingly coherent but actually unreliable information, a phenomenon known as ”hallucination” (Ji et al. 2023a; Cao et al. 2020; Ji et al. 2023b), due to outdated knowledge and a lack of long-tail knowledge (He, Zhang, and Roth 2022; Kandpal et al. 2023; Jiang et al. 2024).
To mitigate the hallucination problem in LLMs, the Retrieval-Augmented Generation (RAG) paradigm has been widely adopted (Izacard et al. 2022; Asai et al. 2023b, a). A common retrieve-and-read framework retrieves relevant information from external knowledge bases, such as Wikipedia, and integrates it with the reader model (i.e., LLMs) to enhance the accuracy of the responses (Gao et al. 2024; Yu et al. 2023a; Lewis et al. 2021). This method leverages rich information from reliable knowledge bases to compensate for the limitations in the internal knowledge of LLMs. However, because the knowledge acquired during LLM pretraining is static, the latest and correct knowledge from external sources can conflict with the parametric knowledge within the model, leading to knowledge conflicts (Xu et al. 2024; Jin et al. 2024). Research indicates that the behavior of LLMs in the face of knowledge conflicts is uncertain, influenced by model parameters and context (Longpre et al. 2022; Xie et al. 2024). In many cases, even when RAG retrieves correct information, LLMs may not necessarily adhere to it when addressing the conflict (Wu, Wu, and Zou 2024). Therefore, enhancing the capability of LLMs to handle knowledge conflicts is a crucial step in improving the performance of RAG.
In response to the aforementioned issue, a mainstream approach is to construct specific instruction-tuning datasets to optimize the knowledge prioritization of LLMs in contexts with varying degrees of relevance (Li et al. 2022; Xue et al. 2023). However, due to the lack of explicit contrastive target training, this prioritization has not been deeply integrated into the model’s self-awareness, resulting in the probability mass still being concentrated on undesirable responses (Tajwar et al. 2024). In real-world scenarios, the context of RAG is complex, and even with relevant contexts, the model may still select irrelevant information from other content to respond  (Shi et al. 2023a; Jiang et al. 2024). Therefore, our insights stem from the error types observed in real-world scenarios involving RAG. We propose introducing negative signals and optimizing knowledge selection in different contexts through preference learning. In the process of learning from human feedback, Direct Preference Optimization (DPO) reparameterizes the reward function to directly fit human preference data. Due to its simplicity and stability, it has been widely adopted and has achieved remarkable results in aligning with human values and intentions. Inspired by this, in this work, we consider utilizing the negative gradient terms in the DPO contrastive objectives to push down the likelihood of undesirable responses in various contexts, further enhancing the adherence capability and noise robustness of LLMs. Although seemingly straightforward, implementing this intuition will encounter the following challenges:
C1. How to more accurately simulate real-world RAG scenarios and introduce more comprehensive and fine-grained negative signals? At the input level, due to factors such as data storage and retriever performance, the context within the retrieval recall window in actual retrieval environments is highly complex (Yao et al. 2024; Gao et al. 2023). At the output level, comprehensive negative signals are crucial for shifting probability mass from undesired responses to desired responses (Tajwar et al. 2024). Therefore, simulating common error scenarios based on complex input context combinations to construct reasonable chosen-rejected answer pairs is a significant challenge.
C2. How to avoid preference imbalance issues in preference optimization? When constructing preference datasets, the length imbalance of chosen-rejected response pairs can lead to length utilization issues, where the fine-tuned model tends to generate longer sentences rather than learning preferences for specific behavior patterns (Park et al. 2024a; Meng, Xia, and Chen 2024). Additionally, the imbalance in preference data representing different behavior patterns (e.g., directly answering ”I don’t know” by ignoring the context versus finding answers from the context) can significantly impact the behavior patterns of the fine-tuned model (Liao et al. 2024). In the process of preference optimization, avoiding behavior pattern imbalances caused by data discrepancies is crucial for simultaneously enhancing the model’s adherence capability and noise robustness.
By jointly considering the above issues, we propose KaPO, a Knowledge-aware Preference Optimization strategy, which constructs comprehensive and balanced preference relations to optimize LLMs’ knowledge selection in different contexts. Specifically, our main contributions are summarized as follows:
  • In the process of constructing the preference dataset, we first simulated real-world RAG scenarios at the input level. We perform refined noise classification based on the relevance between the knowledge context and the question topic, and explore combination methods with evidence to form conflicting context and irrelevant context. At the output level, we simulate two common error types in different context relevance scenarios: Contextual Ignorance and Contextual Overinclusion, and develop training strategies to avoid these errors. This is a general paradigm for constructing knowledge conflict datasets, which can be generalized to various model architectures (Response to Challenge C1).
@@
  • Secondly, we propose a rewriting strategy to address length imbalance and a data ratio balancing strategy to address behavior pattern imbalance, using DPO to optimize LLMs’ adherence capability and noise robustness. These strategies not only eliminate length biases and imbalances in behavior pattern distribution but also enhance the exhaustiveness of the model’s responses. This prevents the degradation of conversational abilities that can occur when training on datasets with shorter answers, such as in reading comprehension tasks (Response to Challenge C2).
@@
  • We validated the training effectiveness of our method on multiple models and datasets and tested its generalization ability in out-of-distribution (OOD) scenarios. The results indicate that our method not only improves the performance of models on test sets but also enhances their adaptability and robustness when confronted with unknown data.
@@ II. Related Work
Numerous studies have explored the behavior of LLMs in knowledge conflict scenarios, providing valuable insights for our work.  Longpre et al. (2022) discovered that large Pre-trained Language Models often prefer to ignore contextual information in favor of the parametric knowledge when facing knowledge conflicts. Recently, with the emergence of LLM such as ChatGPT (OpenAI 2022, 2023) and PaLM (Chowdhery et al. 2022), researchers re-examined this issue.  Wu, Wu, and Zou (2024) highlighted that this tendency to disregard context is strongly influenced by the model’s prior token probability and the deviation degree of the conflicting knowledge. Specifically, conflict knowledge that significantly deviates from the model’s existing knowledge is less likely to be adopted, and knowledge with higher prior probabilities is more challenging to update.  Kassner and Schütze (2019) demonstrated that LLMs are susceptible to being misled by task-irrelevant context. Furthermore,  Tan et al. (2024) indicated that the model’s contextual preferences are closely related to the semantic completeness of the context and its relevance to the question.
Several studies aim to improve the adherence of LLMs to context when faced with knowledge conflicts. For instance, Knowledge Aware Fine-Tuning (KAFT) (Li et al. 2022) attempts to enhance models’ ability to utilize external knowledge by constructing challenging counterfactual knowledge from standard training datasets, as well as incorporating irrelevant knowledge to improve models’ noise resistance. However, as previously mentioned, the applicability of this approach in real-world RAG scenarios is limited. Additionally, decoding-based methods (Jin et al. 2024; Chen, Zhang, and Choi 2022), such as Context-Aware Decoding (CAD) (Shi et al. 2023b), adjust the output probabilities of LLMs during token generation in a manner similar to contrastive decoding, conditioned on the relevant context. However, this approach may impact the semantic coherence of long responses. Moreover, prompt-based methods employ sophisticated designed prompts to ensure that LLMs adhere to the provided context (Si et al. 2023; Zhou et al. 2023). Nevertheless, related research indicates that merely modifying prompts does not significantly alter LLM’s internal prior token probabilities (Wu, Wu, and Zou 2024), potentially limiting the effectiveness of this approach.
@@
RAG incorporates the external knowledge retrieval component via prompt engineering to achieve more factual consistency, enhancing the reliability and interpretability of LLMs’ responses (Lewis et al. 2021). Some studies have made improvements to the retrieve-and-read framework by generating intermediate contexts using the parameter knowledge acquired during the pretraining phase, thereby enhancing the quality of the final response. These intermediate contexts may include commonsense knowledge (Liu et al. 2022), domain-specific knowledge (Luo et al. 2023), and chain-of-thought(COT) reasoning process (Wei et al. 2023; Kojima et al. 2023; Li et al. 2023). Furthermore,  Zhao et al. (2023a),  Wang et al. (2023a) and  Yu et al. (2023b) have utilized retrieved knowledge to edit the intermediate contexts of the COT process, thereby updating conflicting knowledge. However, these intermediate contexts generated by LLMs may contain hallucinations or other inaccurate information, potentially misleading the retrieval or reader models. Additionally, the frequent interactions with LLMs result in inefficiencies in real-world applications (Jiang et al. 2024).
@@ A. Task Definition
Given an LLM $\Theta$ and an input natural language question $q$, we ask $\Theta$ to generate a response $\alpha=\Theta(q)$, which represents the parametric knowledge for $q$. Besides, a typical retrieve-and-read framework can be expressed as $y=\Theta(q|\tau)$, where $y$ is the output of $\Theta$ based on $\tau$. Context $\tau$ is a permutation of $D_{j}^{r},j=1,2,\ldots,K$, which represents a set of documents retrieved based on $q$. Assume $S=\{a_{i}\},i=1,2,\ldots,N$ constitutes the set of correct answer, each of which is derived from the retrieved document $D_{i}^{r}$. Note that $K$ is not necessarily equals with $N$, because some retrieved documents may not contain any answer for $q$ which are known as noises.
It’s clearly that $\alpha$ and $a_{i}$ are independent. Knowledge conflict appears when $\alpha\notin S$, and at this time response $y$ of $\Theta(q|\tau)$ can be uncertain. To simplify the discussion, we limit $N$ to a maximum of $1$, which means context $\tau$ contains at most one document $D_{1}^{r}$ from which the answer can be derived. Our purpose is to make sure $y=a_{1}$ when $|S|=1$ and $y=\alpha$ when $|S|=0$. In other word, LLM $\Theta$ should use appropriate external knowledge when there exists a document which contains the necessary knowledge regardless of conflicting with parameter knowledge, while use its parameter knowledge when retrieved documents are all irrelevant.
@@ B. Contradictory Knowledge
Constructing knowledge that conflicts with LLM’s parameter knowledge is crucial to condition $|S|=1$. For question $q$ in RAG scenarios, this conflict is reflected in conflicting answers $a_{i}$ which are inconsistent with LLM’s parameter answer $\alpha$. It is important to note that these conflicting answers $a_{i}$ do not necessarily have to be correct, nor is the LLM’s parameter answer $\alpha$ always incorrect. In our approach, both answers can be incorrect to the question as long as they conflict with each other. The key to knowledge conflict lies in the conflict itself, regardless of correctness. This addresses a common misconception in previous work, where researchers often ensured that one answer was correct and the other incorrect (Tan et al. 2024; Wu, Wu, and Zou 2024), which not only increased the difficulty of data filtering but also overlooked some knowledge conflict scenarios.
Specifically, we first extract world knowledge acquired during the pretraining phase of the large model, marked as parameter answer $\alpha$. We encourage LLM to abstain from answering when uncertain and normalize all instances of refusal. Additionally, we refine the response formats for other parametric knowledge. The revised results are presented in Table 1.
For a given question $q$ and LLM’s parameter answer $\alpha$, there are two potential sources of conflicting answers. The first source is the realistic answer $a_{r}$ to the question. The second source is a fabricated answer $a_{c}$ generated using GPT-4 that deviates from the realistic answer $a_{r}$. The latter is often referred to as a counterfactual answer, which we require to be as plausible as possible. Thus, for a question $q$ and LLM’s parameter answer $\alpha$, we can obtain at least one conflicting answer, ensuring that this answer is not excessively far-fetched. The few-shot prompts are structured as follows:
Prompt 3.1: Extract Parameter Answer This is a question about {Title}. Please answer the question {Question $q$}. Please provide a direct answer without analysis. If you are unsure or do not know the answer, please respond with ‘I don’t know’.
Prompt 3.2: Generate Counterfactual Answer Please generate speciously plausible but incorrect answer to the question. Provide only the false answers; do not reiterate the queries. Question: What is the capital of France? Answer: Paris. Fake answer: Lyon. Question: What is the highest mountain in the world? Answer: Mount Everest. Fake answer: Lhotse. …7 more examples … Question: Who is the founder of Microsoft? Answer: Bill Gates. Fake answer: Steve Jobs. Question: {Question $q$} Answer: {Realistic Answer $a_{r}$} Fake answer:

### v2

# KnowPO: Knowledge-aware Preference Optimization for Controllable Knowledge Selection in Retrieval-Augmented Language Models

@@ Abstract
By integrating external knowledge, Retrieval-Augmented Generation (RAG) has become an effective strategy for mitigating the hallucination problems that large language models (LLMs) encounter when dealing with knowledge-intensive tasks. However, in the process of integrating external non-parametric supporting evidence with internal parametric knowledge, inevitable knowledge conflicts may arise, leading to confusion in the model’s responses. To enhance the knowledge selection of LLMs in various contexts, some research has focused on refining their behavior patterns through instruction-tuning. Nonetheless, due to the absence of explicit negative signals and comparative objectives, models fine-tuned in this manner may still exhibit undesirable behaviors such as contextual ignorance and contextual overinclusion. To this end, we propose a Knowledge-aware Preference Optimization strategy, dubbed KnowPO, aimed at achieving adaptive knowledge selection based on contextual relevance in real retrieval scenarios. Concretely, we proposed a general paradigm for constructing knowledge conflict datasets, which comprehensively cover various error types and learn how to avoid these negative signals through preference optimization methods. Simultaneously, we proposed a rewriting strategy and data ratio optimization strategy to address preference imbalances. Experimental results show that KnowPO outperforms previous methods for handling knowledge conflicts by over 37%, while also exhibiting robust generalization across various out-of-distribution datasets.
@@ I. Introduction
Large Language Models (LLMs) (Taylor et al. 2022; Zhao et al. 2023c) have been widely applied in various fields, such as natural language processing, question-answering systems, and text generation, giving rise to numerous AI applications (Kaplan et al. 2020; Vu et al. 2024). These models exhibit outstanding performance in many tasks, primarily due to their large-scale parameters and extensive pre-training data (Ziegler et al. 2020; Wang et al. 2023d). However, because of the static nature of the training data, LLMs may generate seemingly coherent but actually unreliable information, a phenomenon known as “hallucination” (Ji et al. 2023a; Cao et al. 2020; Ji et al. 2023b), due to outdated knowledge and long-tail knowledge (He, Zhang, and Roth 2022; Kandpal et al. 2023; Jiang et al. 2024). Retrieval-Augmented Generation (RAG) paradigm (Izacard et al. 2022; Asai et al. 2023b, a), within a retrieve-and-read framework, leverages information from reliable knowledge bases to compensate the static nature of the Internal knowledge of LLM. However, the performance of RAG framework is limited by the knowledge conflicts between internal knowledge stored in LLM parameters and external database (Xu et al. 2024; Jin et al. 2024). In this paper, we focus on resolving knowledge conflict by adhere to the retrieved knowledge, and meanwhile, improving the robustness against noise in the retrieved context.
In response to the aforementioned issue, a mainstream approach is to construct specific instruction-tuning datasets to optimize the knowledge prioritization of LLMs in contexts with varying degrees of relevance (Li et al. 2022; Xue et al. 2023). However, as shown in Figure 1, achieving a balance between adherence capability and noise robustness is highly challenging. On one hand, when the LLM heavily relies on external knowledge, it risks over-focusing on irrelevant retrieval contexts, struggling to effectively discern noise. On the other hand, an excessive emphasis on enhancing the LLM’s noise resistance can inadvertently filter out useful contextual information (Wu, Wu, and Zou 2024). Moreover, the manifestation of these capabilities is closely related to the complexity of the context in real-world RAG scenarios (Longpre et al. 2022; Xie et al. 2024). Therefore, it is crucial to address the balance between adherence capability and noise robustness in real RAG scenarios.
Figure 1: An illustrative example of how does LLM behave when encountering knowledge conflicts in RAG scenarios.
Therefore, our insights stem from the error types observed in real-world scenarios involving RAG. We observe that existing research fails to distinguish between supervisory signals, leading to adherence capability and noise robustness being treated as analogous instruction-following positive examples. This leads to contradictory signals during instruction-following training, causing learning variance and impeding the effective acquisition of both capabilities. To address this, we propose a more nuanced optimization approach that introduces preference data specifically describing adherence capability and noise robustness. Leveraging efficacious and extensively utilized Direct Preference Optimization (DPO) (Rafailov et al. 2024), we optimize the model’s ability to leverage external knowledge, thereby enhancing the overall efficacy of RAG.
Although seemingly straightforward, implementing this intuition faces these challenges: (C1) How to more accurately simulate complicated context in real-world RAG scenarios and introduce more comprehensive, fine-grained negative signals? (C2) How to resolve data discrepancies in preference learning to avoid behavior pattern imbalances?
@@
By jointly considering the above issues, we propose KnowPO, a Knowledge-aware Preference Optimization strategy, which constructs comprehensive and balanced preference relations to optimize LLMs’ knowledge selection in different contexts. i) Specifically, we simulated real-world RAG scenarios at the input level. We perform refined noise classification based on the relevance between the knowledge context and the question topic, and explore combination methods with evidence to form conflicting context and irrelevant context. At the output level, we simulate two common error types in different context relevance scenarios: Contextual Ignorance and Contextual Overinclusion, and develop training strategies to avoid these errors. ii) Secondly, we propose a rewriting strategy to address length imbalance and a data ratio balancing strategy to address behavior pattern imbalance, using DPO to optimize LLMs’ adherence capability and noise robustness. These strategies not only eliminate length biases and imbalances in behavior pattern distribution but also enhance the exhaustiveness of the model’s responses. This prevents degradation of conversational abilities that can occur when training on datasets with shorter answers, such as in reading comprehension tasks.
Our main contributions are summarized as follows:
  • We observed that LLMs in RAG scenarios fail to effectively balance adherence capability and noise robustness. and thus proposed the KnowPO framework, which refines negative supervisory signals to enhance LLM behavior when encountering knowledge conflicts.
  • We proposed a general paradigm for constructing knowledge conflict datasets, comprehensively covering various error types and generalizable to different model architectures. We also proposed a rewriting strategy and data ratio optimization strategy to address preference imbalances.
@@
  • We validated our method’s training effectiveness on multiple models and datasets and tested its generalization ability in out-of-distribution (OOD) scenarios. The results indicate that our method not only improves the performance of models on test sets but also enhances their adaptability and robustness when confronted with unknown data.
@@ II. Related Work
Numerous studies have explored LLMs’ behavior in knowledge conflict scenarios, providing valuable insights for our work.  Longpre et al. (2022) discovered that large Pre-trained Language Models often prefer parametric knowledge over contextual information when facing knowledge conflicts. Recently, with the emergence of LLM such as ChatGPT (OpenAI 2023) and PaLM (Chowdhery et al. 2022), researchers re-examined this issue.  Wu, Wu, and Zou (2024) highlighted that this tendency to disregard context is influenced by the model’s prior token probability. Specifically, parametric knowledge with higher prior probabilities is more challenging to update to conflicting knowledge.  Kassner and Schütze (2019) demonstrated that LLMs are susceptible to being misled by task-irrelevant context. Furthermore,  Tan et al. (2024) indicated that the model’s contextual preferences are linked to the semantic completeness of the context and its relevance to the question.
Several studies aim to improve the adherence of LLMs to context amid knowledge conflicts. For instance, Knowledge Aware Fine-Tuning (KAFT) (Li et al. 2022) enhances models’ ability to use external knowledge by creating challenging counterfactual knowledge from training datasets and incorporating irrelevant knowledge to boost noise resistance. However, as previously mentioned, the applicability of this approach in real-world RAG scenarios is limited. Additionally, decoding-based methods (Jin et al. 2024; Chen, Zhang, and Choi 2022), like Context-Aware Decoding (CAD) (Shi et al. 2023b), adjust LLMs’ output probabilities during token generation, akin to contrastive decoding, conditioned on relevant context. However, this approach may impact the semantic coherence of long responses. Moreover, prompt-based methods employ sophisticated designed prompts to ensure that LLMs adhere to the provided context (Si et al. 2023; Zhou et al. 2023a). However, research shows that merely modifying prompts doesn’t significantly alter LLMs’ internal prior token probabilities (Wu, Wu, and Zou 2024), potentially limiting the effectiveness of this approach.
@@
RAG incorporates the external knowledge retrieval component via prompt engineering to achieve more factual consistency, enhancing the reliability and interpretability of LLMs’ responses (Gao et al. 2024; Yu et al. 2023a; Lewis et al. 2021). Some studies have made improvements to the retrieve-and-read framework by generating intermediate contexts using the parameter knowledge acquired during the pretraining phase, thereby enhancing the quality of the final response. These intermediate contexts may include commonsense knowledge (Liu et al. 2022), domain-specific knowledge (Luo et al. 2023), and chain-of-thought(COT) reasoning process (Wei et al. 2023; Kojima et al. 2023; Li et al. 2023). Furthermore,  Zhao et al. (2023a),  Wang et al. (2023a) and  Yu et al. (2023b) have utilized retrieved knowledge to edit the intermediate contexts of the COT process, thereby updating conflicting knowledge. However, these intermediate contexts generated by LLMs may contain hallucinations, potentially misleading the retrieval or reader models. Additionally, the frequent interactions with LLMs result in inefficiencies in real-world applications (Jiang et al. 2024).
@@ A. Task Definition
Given an LLM $\Theta$ and an input natural language question $q$, we ask $\Theta$ to generate a response $\alpha=\Theta(q)$, representing the parametric knowledge for $q$. Assume in a typical retrieve-and-read framework, context $\tau$ is a permutation of $D_{j}^{r},j=1,2,\ldots,K$, which represents a set of documents retrieved based on $q$. And $S=\{a_{i}\},i=1,2,\ldots,N$ constitutes the set of contextual answer, each of which is derived from a retrieved document $D_{\tau_{i}}^{r}$. We can simplify the RAG task into $y=\Theta(q\|\tau)$, where $y$ is output of $\Theta$ based on context $\tau$. Note that $K$ is not necessarily equals with $N$, because some retrieved documents may not contain any answer for $q$ and are known as noises.
It’s clearly that $\alpha$ and $a_{i}$ are independent. Knowledge conflict appears when $\alpha\notin S$, and at this time response $y$ of $\Theta(q\|\tau)$ can be uncertain. To simplify the discussion, we limit $N$ to a maximum of $1$, which means context $\tau$ contains at most one document $D_{\epsilon}^{r}$ from which the answer can be derived. Our purpose is to make sure $y=a_{\epsilon}$ when $|S|=1$ and $y=\alpha$ when $|S|=0$. In other word, LLM $\Theta$ should use appropriate external knowledge when there exists a document which contains the necessary knowledge regardless of conflicting with parameter knowledge, while use its parameter knowledge when retrieved documents are all irrelevant.
@@ B. Contradictory Knowledge
Constructing knowledge that conflicts with LLM’s parameter knowledge is crucial to condition $|S|=1$. For question $q$ in RAG scenarios, this conflict is reflected in conflicting answers $a_{cf}$ which are inconsistent with LLM’s parameter answer $\alpha$. It is important to note that these conflicting answers $a_{cf}$ do not necessarily have to be correct, nor is the LLM’s parameter answer $\alpha$ always incorrect. In our approach, both answers can be incorrect to the question as long as they conflict with each other. The key to knowledge conflict lies in the conflict itself, regardless of correctness. This addresses a common misconception in previous work, where researchers often ensured that one answer was correct and the other incorrect (Tan et al. 2024; Wu, Wu, and Zou 2024), which not only increased the difficulty of data filtering but also overlooked some knowledge conflict scenarios.
Specifically, we first extract world knowledge acquired during the pretraining phase of the large model, marked as parameter answer $\alpha$. We encourage LLM to abstain from answering when uncertain. Additionally, we refine the response formats for other parametric knowledge. The revised results are presented in Table 1.
For a given question $q$ and LLM’s parameter answer $\alpha$, there are two potential sources of conflicting answers $a_{cf}$. The first is the realistic answer $a_{real}$ to the question. The second is a fabricated answer $a_{ctf}$ generated using GPT-4 that deviates from the realistic answer $a_{real}$. The latter is often referred to as a counterfactual answer, which we require to be as plausible as possible. Thus, for a question $q$ and LLM’s parameter answer $\alpha$, we can obtain at least one conflicting answer, ensuring it is not overly far-fetched. The few-shot prompts can be found in Appendix A.